key.sty

%%
%% This is file `key.sty', generated
%% on <1995/10/11> with the docstrip utility (2.2i).
%%
%% The original source files were:
%%
%% key.dtx  (with options: `package')
%% Package to draw keys in boxes with rounded corners.
\ProvidesPackage{key}[95/10/11 ver 2.0a Creates a boxed key image]
\RequirePackage{fancybox}
\newcommand{\key}[1]%
{%
   \Ovalbox%
   {%
      \makebox[0.7em]%
      {%
         {\small\textsf{#1}}%
         \rule[-0.0em]{0pt}{0.7em}%
      }%
   }%
}
\endinput
%%
%% End of file `key.sty'.